OpenStudy (anonymous):

Need help, Calculate the wavelength of a proton with energy 2.45eV.

OpenStudy (anonymous):

I cant get the write answer in meter!!! Calculate the wavelength of a proton with energy 2.45eV.

OpenStudy (aaronq):

did you use the deBroglie equation?

OpenStudy (aaronq):

theres one specific for things moving really fast (near c) and ones that aren't

OpenStudy (anonymous):

Yes but nothing good. I came up with 5.05401530612e-07 but it's not right. Would you please check it for me...

OpenStudy (aaronq):

how did you do it?

OpenStudy (anonymous):

E=hc/lamb 1ev=1.6*10^-19 Wavelength=hc/E

OpenStudy (aaronq):

the first equation is for photons, you need to use use: KE=mv^2 to find v, velocity then, the deBroglie eqn: lambda = h/mv

OpenStudy (anonymous):

Many thanks... It worked out ;)

OpenStudy (aaronq):

good stuff, no prob !